### v1

# XpertAI: uncovering model strategies for sub-manifolds

Abstract
In recent years, Explainable AI (XAI) methods have facilitated profound validation and knowledge extraction from ML models. While extensively studied for classification, few XAI solutions have addressed the challenges specific to regression models. In regression, explanations need to be precisely formulated to address specific user queries (e.g. distinguishing between ‘why is the output above 0?’ and ‘why is the output above 50?’). They should furthermore reflect the model’s behavior on the relevant data sub-manifold. In this paper, we introduce XpertAI, a framework that disentangles the prediction strategy into multiple range-specific sub-strategies and allows formulating precise queries about the model (the ‘explanandum’) as a linear combination of those sub-strategies. XpertAI is formulated generally to work alongside popular XAI attribution techniques, based on occlusion, gradient integration, or reverse propagation. Qualitative and quantitative results, demonstrate the benefits of our approach.
Machine Learning, ICML
Figure 1: Left: Methodological overview of our proposed XpertAI approach. We introduce a virtual layer of range experts to get insights into the model behavior on output-range specific sub-manifolds. In the forward pass, we retain the model behavior through the additive structure of experts (top, right). At the same time, they represent a well-organized latent space that can be systematically queried when explaining the model (bottom left). Right: While the naive application of attribution methods typically answers questions from a generic point of view (grey) our approach enables answers to more nuanced queries as defined by the user (red). For the task of wine quality prediction, we see significant structural changes in the explanation for different reference values (center). When analyzing the losses of a wind turbine, contextualized attributions correspond more faithfully to the true losses (right).
1 Introduction
Machine learning has provided powerful predictive models for numerous scientific and industrial applications. As the use of ML models for critical autonomous decisions increases, there is a growing demand for establishing trust while maintaining their predictive capabilities. Explainable artificial intelligence (XAI) has emerged as a step towards enhancing transparency and allows for insights into the inner workings of these highly complex AI models (Arrieta et al., 2020; Samek et al., 2021). XAI can be utilized for both, model validation against expert intuition and for obtaining new insights into the data-generating processes under investigation (Krenn et al., 2022; Klauschen et al., 2024).
So far, the predominant focus within XAI has been placed on understanding the decisions made by classification models (Baehrens et al., 2010; Zeiler & Fergus, 2014; Bach et al., 2015; Zhou et al., 2016; Ribeiro et al., 2016). The widely used family of post-hoc attribution methods aims to achieve this by allocating evidence for a particular class across the corresponding input features. Thus they indicate the extent to which each feature has contributed to the model output. In this process, the model’s decision boundary serves as a natural point of reference for the explanation.
In regression, on the other hand, the matter is more complex. There, the equivalent to the decision boundary needs to be defined for every single query, since it is a priori unknown which of the two questions ‘why is the output above 50?’ or ‘why is the output above 0?’ is most relevant for the user (Letzgus et al., 2022). Moreover, in non-linear problems, sub-manifolds on which the model builds specific responses are to be expected, for example, for different output values: Consider, for example, the task of predicting the quality of wine based on some descriptive features (year, grape, price, etc.). Intuitively, the explanation of what makes a premium wine better than a decent alternative is likely to differ significantly from the explanation of what makes it better than a low-quality rotgut, which should be reflected in the attributions.
To address these challenges, we propose our XpertAI framework. The basic idea is to decompose the regression model output into a set of additive basis functions, so-called range experts (compare Figure 1). Each range expert is dedicated to capturing the model behavior within a specific, output-range-dependent sub-manifold. Subsequently, range experts can be queried by the user with any state-of-the-art attribution method to obtain explanations that are contextualized to the individual explanatory needs. We demonstrate the benefits of our method on several (controlled and real-world) problems (see Figure 1). Regarding the wine example from above, we indeed find that the model considered the price the most important input feature to distinguish an excellent wine from a bad one. But when explaining with respect to decent alternatives (close-by-reference values), other quality-related features become much more important. In another case study, we use attributions to monitor the performance of a wind turbine. There, we find that our contextualized explanations more faithfully capture the performance losses, which enables better maintenance decisions in practice. In addition to these qualitative insights, we report improved faithfulness through better contextualization with XpertAI.
2 Related Work
Our proposed method relates to several specific areas of XAI, which we will briefly discuss below (see e.g. (Samek et al., 2021; Arrieta et al., 2020) for XAI reviews).
2.1 Mixture of Experts
The concept of Mixture of experts (MoE) has been used for several decades for the purpose of learning better models (Jacobs et al., 1991; Hampshire & Waibel, 1992; Pawelzik et al., 1996). Beyond model improvement, recent work has demonstrated the value of the MoE approach for achieving more transparent models by combining inherently interpretable linear experts (Ismail et al., 2023). In contrast, our approach utilizes MoEs for explaining models in a post-hoc manner, without restriction on the structure of the model, and steering the expert to become ‘range experts’ focusing on specific value ranges.
2.2 Context in XAI attribution methods
Generally speaking, every explanation requires context to be meaningful. This also applies to XAI attribution methods which incorporate a baseline in input space $\widetilde{x}$ or latent space $\widetilde{a}$ to do so. Occlusion-based methods, such as Shapley values (Lundberg & Lee, 2017), for example, need the user to select a baseline in input space to occlude with. Gradient Integration methods (Sundararajan et al., 2017) need to define a start of the integration path and deep Taylor expansions (Montavon et al., 2017) root points. Propagation-based methods allow for a baseline selection in latent-space, either explicitly (Shrikumar et al., 2017) or implicitly via the choice of propagation rules (Montavon et al., 2017; Bach et al., 2015). If we have a concrete counterfactual sample $x^{\prime}$ it can therefore be utilized to obtain a contextualized result (Albini et al., 2022; Dombrowski et al., 2023). In practice, however, this is often not the case which leaves the selection to the user.
Finding an appropriate baseline or root point is a difficult task. At the same time, it has considerable influence on the explanation (Sundararajan & Najmi, 2020). Suitable candidates have to be on the data manifold, preferably in proximity to the sample in question (Dombrowski et al., 2023). In classification, a baseline additionally has to lay on the decision boundary to obtain meaningful attributions. In regression, each $\widetilde{x}$ inherently corresponds to a reference value $f(\widetilde{x})=\widetilde{y}$ on the model output scale, which adds yet another degree of freedom to be considered.
One solution that comes to mind is to compute expected attributions across multiple baselines (Kapishnikov et al., 2019; Lundberg & Lee, 2017). For example, by conditionally drawing from the dataset. The conditional mean around some desired reference value ($x^{\prime}=\mathbb{E}(x|\widetilde{y})$) could also serve as a synthetic counterfactual baseline, although $f(x^{\prime})$ usually does not coincide with the exact desired $\widetilde{y}$ for non-linear models. Alternatively, Bayesian- or generative modelling offer a computationally more expensive way to synthesize a counterfactual (Van Looveren et al., 2021; Spooner et al., 2021). In this contribution, we propose a practical solution that ensures contextualization by design and therefore increases robustness against suboptimal baselines in input space.
2.3 Disentangled XAI and Virtual Layers
While refining the question to be asked is essential in a regression setting (see previous sections), many works have focused on independently refining the explanation itself (mainly in a classification context). Specifically, enriching explanations by identifying its multiple components, associated with distinct abstract concepts. These can be obtained in a supervised manner (Kim et al., 2018; Zhao et al., 2023), in an unsupervised manner (Vielhaben et al., 2023; Chormai et al., 2022), or by directly inspecting neurons (Zeiler & Fergus, 2014; Zhou et al., 2016; Achtibat et al., 2023).
This kind of analysis often involves an informed transformation of latent representations to obtain a meaningful or relevant ’concept space’, followed by the inverse transformation to leave the overall model behaviour intact. Therefore, these approaches are referred to as virtual layers. (Chormai et al., 2022), for example, extract sub-concepts that jointly contribute to the explanation of an overall class concept. Likewise, (Vielhaben et al., 2024) generates a Fourier basis on which the prediction of speech samples can be more efficiently analyzed, and (Linhardt et al., 2024) introduces a virtual PCA layer, which disentangles verified from unverified factors of variation and subsequently prune the latter for increased robustness.
We extend these prior efforts to the broad domain of regression, by introducing a novel technique aiming at disentangling global phenomena that exert influence consistently across the entire range of potential regression outputs from more localized, context-specific patterns (see Figure 1).
Figure 2: Diagram of our two-step approach for obtaining fine-grained explanations from an existing regression model. The first step consists of adding a collection of range experts to the model. The second step synthesizes a query $q$ from those range experts and produces a corresponding explanation (the exemplary query on the right is sigmoidal with the ML model’s output but linear with the experts).
3 Our Method: XpertAI
In the following, we introduce our novel method, called XpertAI for explaining neural network regression models. Our approach is inspired by the MoE concept and consists of appending range experts to a given ML model, thus allowing the user to formulate precise queries for which range they need an explanation. This appendage can be seen as a virtual layer inserted in the neural network, which – while leaving the overall prediction function intact – enriches it by providing the basis for query formulation and explanation.
3.1 Adding Range Experts
We abstract the ML model as a function $f$ mapping the input $x$ to a real-valued output $y$. The model may either be a pure black-box or a neural network with multiple layers. We define the range experts as the following collection of functions building on the output of the ML model:
$z=\left(\begin{array}[]{l}\rho_{0,\tau}(y)\\ \rho_{0,\tau}(y-\tau)\\ \rho_{0,\tau}(y-2\tau)\\ \vdots\end{array}\right)$ (5)
where $\rho_{0,\tau}(y)=\min(\max(y,0),\tau)$ clips the input to the interval $[0,\tau]$. A low $\tau$ corresponds to a more specialized expert. The transformation in Eq. (5) is also known as thermometer coding. The architecture that results from appending these experts is shown in Fig. 2. Assuming the values of $y$ are always positive (which we can ensure through offsetting) we can reconstitute the output prediction by summing the experts’ outputs:
$\textstyle y=\sum_{m}z_{m}$ (6)
The mapping from $y$ to $z$ and back to $y$ can be seen as a virtual layer whose insertion does not affect the input-output mapping, but that provides additional functionality. Unlike previous formulations of virtual layers (Vielhaben et al., 2024), ours is placed at the output, enabling a disentanglement of the explanation in terms of those range experts.
Consider now the task of attribution. Classical explanation techniques would attribute $y$ to the features of $x$ (something we denote by $\mathcal{E}(y,x)$). The virtual layer allows us to compose two attribution steps:
$R_{m}$ $=\mathcal{E}(y,z)_{m}$
$R_{im}$ $=\mathcal{E}(R_{m},x)_{i}$
where $R_{m}$ denotes the contribution of expert $m$ to the output $y$ (in our case we simply have $R_{m}=z_{m}$), and $R_{im}$ can be interpreted as the contribution of input feature $i$ through expert $m$. The overall explanation can be seen as a matrix of size # features $\cdot$ # experts, from which it will be possible to formulate and answer precise user queries.
3.2 Querying / Explaining
Whereas the disentanglement performed above provides a more detailed view of the prediction behavior than a simple explanation, the user is often interested in particular aspects of it. Our approach lets the user formulate a query (or ‘explanandum’) as a linear combination of the range experts:
$\textstyle q=\sum_{m}w_{m}z_{m}$ (7)
An example of such a query is given in Fig. 2 (right). For example, if the user is interested in what makes a prediction $y=60$ larger than a reference value of $50$, the query $q$ can be shaped in the form of a sigmoid centered at the reference value $50$.
Once a query has been prepared (i.e. once the weights $w_{m}$ have been defined), an explanation to that query $\mathcal{E}(q,x)$ can be generated by any state-of-the-art attribution method:
$\textstyle\mathcal{E}(q,x)=\mathcal{E}\big{(}\sum_{m}w_{m}z_{m},x\big{)}$ (8)
Note that for explanation techniques that fulfill the linearity axiom w.r.t. the last layer of representation, we can further develop the expression of the explanation as:
$\mathcal{E}(q,x)=\sum_{m}w_{m}\mathcal{E}\big{(}z_{m},x\big{)}$ (9)
It shows that the explanation is a linear combination of the explanations of all basis elements $z_{m}$. This formulation can be advantageous when the explanation is associated with many different queries or when the query arrives in real-time, in which case the explanation basis can be pre-computed. We note that our approach satisfies some key desirable properties of an explanation:
Proposition 1 (Conservation).
If $\forall_{m}:\sum_{i}\mathcal{E}(z_{m},x)_{i}=z_{m}$, then $\sum_{i}\mathcal{E}(q,x)_{i}=q$, in other words, if each range expert $z_{m}$ can be attributed to input features in a conservative manner, then explanations of any query $q$ are also conservative.
Proposition 2 (Irrelevance).
If $\forall_{m}:\mathcal{E}(z_{m},x)_{i}=0$, then $\mathcal{E}(q,x)_{i}=0$, in other words, if we verify that for a given data point, the feature is irrelevant for all range experts, then it is also irrelevant for any query built on those experts.
These two results are easily retrievable by observing the specific structure of the explanation given in Eq. (9). Proofs can be found in Appendix A.
3.3 Structural Disentanglement
When the underlying explanation method relies not directly on the ML model’s output but on its computational graph (e.g. LRP), the latter must be disentangled. Clearly, such a structural disentanglement is missing as the mapping from activations $a$ to the expert’s outputs $z$ passes through a one-dimensional bottleneck $y$ (the original real-valued output). We propose to replace the original mapping $a\mapsto(z_{m})_{m}$ by a surrogate model
[…]
where the second part of the mapping is given by $\widehat{z}_{m}=\rho_{0,\tau}(s_{m})$, a hard-coded saturation forcing the surrogate and true experts to produce outputs in the same range. We then build for each expert the loss function:
$\ell(s_{m},z_{m})=\left\{\begin{array}[]{ll}\max(0,s_{m})&z_{m}\leq 0\\[1.42262pt] |s_{m}-z_{m}|&0<z_{m}<\tau\\[1.42262pt] \max(0,\tau-s_{m})&z_{m}\geq\tau\end{array}\right.$
which encourages that the surrogate’s output is correct within-range and on the correct side outside-range. We then solve $\min_{\theta}\mathbb{E}[\sum_{m}\ell(s_{m},z_{m})]$ with $\mathbb{E}[\cdot]$ denoting the expectation over the training data. To preserve not only the prediction output of the original model but also its prediction strategy (i.e. the feature it uses) further steps are needed. One approach is to enforce the loss function not only on the data but also on perturbations of the data (Stutz et al., 2019). For example, activations can be randomly turned off (with a probability chosen between 00 and $1$). This perturbation scheme ensures in particular that the Shapley value explanations of the original and disentangled models become similar (i.e. that they predict the same for the same reasons). Furthermore, we find that freezing the bias in the output layer is important in order to achieve the desired structural disentanglement.
4 XpertAI-Opinion: insights into model behavior on sub-manifolds
We now demonstrate how our XpertAI approach can help users disentangle local and global effects for meaningful insights in three different case studies. For each of them, we briefly introduce the dataset, model and XpertAI setting, before presenting the insights. We present results from using both, Integrated Gradients and LRP. Details on all case studies can be found in Appendix C.
4.1 Uncovering output-scale-specific strategies
We start this investigation in a rather controlled environment. For that, we adopt the world’s arguably best-known classification data set, MNIST (Deng, 2012), and transform it into a regression problem (rMNIST). For simplicity, we take the subset of only three digits (4,5, and 6) and calculate a one-dimensional t-SNE representation (van der Maaten & Hinton, 2008), which henceforth serves as a new label for each sample. Additionally, we ensure labels are distributed uniformly between values of zero and one. As a result, the individual digits populate continuous parts of the output dimension (in our case sorted by digit magnitude, which facilitates interpretation) while sorting within each digit bin is based on the respective digit’s rotation (compare Figure 3). We now train a vanilla CNN model architecture to learn this mapping from image to output scale. For contextualized insights, we train three range experts (one for each digit range). We first discuss qualitative results and present its quantitative evaluation in Section 5.
Figure 3: Examples from three classes of the MNIST dataset for handwritten digit recognition (left) mapped to a real-valued scale with the help of a one-dimensional t-SNE embedding (right). Digits populate continuous ranges of the new target, and sorting within the digit ranges corresponds to digit rotations.
Figure 4 shows a comparison between the standard and the XpertAI explanations for both, IG and LRP. In the bottom row, we show the average naive attributions over all samples within the respective output range. The explanations from the lowest-range expert (second row from below) are very similar to the original ones since they implicitly explain relative to the same reference value. The range experts of the upper digit ranges (marked in colors), however, enable more granular insights. It is visible how the range experts for the output range populated with the digits five and six (marked in orange and blue) seem to focus specifically on the rotation of the digit. A rotation to the right is associated with lower values (negative attribution, blue) and vice versa. See Appendix C for more basis functions. In Section 5, we will see that these qualitative differences in attributions also result in improved quantitative evaluation scores for attribution faithfulness.
Figure 4: Mean attributions over different cohorts of samples (columns) and basis functions (rows). The bottom row represents naive attributions. The top three rows correspond to ascending range-specific XpertAI bases. Note, how only the latter exposes the digit rotation within the respective digit ranges (orange/blue).
4.2 What sets apart a good wine from an excellent one?
As mentioned in the introduction, we now turn to a slightly more hedonistic and tangible example where many people have strong opinions: the quality of red wine. We utilize Kaggle’s Spanish red wine dataset [^1] https://www.kaggle.com/datasets/fedesoriano/spanish-wine-quality-dataset which contains several thousand wine samples. They are described by five numerical (year, price, as well as body, acidity, and quality scores) and four categorical (name of the winery and the wine, grape, region) features. As a regression target, we have selected the quality score which is an ’average rating’ given by thousands of testers (rating binned into 8 discrete quality levels), and we utilize the remaining eight features as inputs. After data-pre-processing around 1700 samples are left. We have trained a small fully-connected ANN which achieved an $R^{2}$ of around 0.7.
We now want to learn what, according to the model, sets apart a good wine from an excellent one. We define wine as good when it belongs to the top 10 % and excellent when it belongs to the top 1.5 % of the model output range. We train three range experts ($\tau=0.33$) and compare the respective attributions obtained from standard IG with its application within the XpertAI framework. Figure 5 shows the decomposition of the excellent wine attributions into the respective expert bases. Aside from the natural change in attribution scales, the most prominent difference in the explanations is the contribution of the price to the model outcome. For the naive IG attributions, the price is the by far most important feature (meaning high prices alone are the main indicator for excellent wines). The contextualized XpertAI attributions, on the other hand, give a much more balanced picture. Here, the outcome suggests that the price is the most important feature only for the low-quality range expert (meaning what distinguishes an average from a poor wine, blue). The relative importance of the price, however, is significantly reduced when compared to average wines (orange) and almost vanishes when compared to good wines (greed). There, the sum of all other quality criteria is much more important than the price of the wine itself. This directly translates to some actionable (and intuitive) insight: if you next time buy a wine in the supermarket, don’t go cheap to ensure you buy a decent wine. When looking for an excellent one though, you might be better off with the expert judgement of your local wine seller.
Figure 5: Decomposition of naive explanations (grey) for samples from the high output range (’excellent’ wines) with respect to low, medium and high-quality reference values (colorful). The XpertAI explanations allow for nuanced insights into what makes an excellent wine better than the worst (blue), a decent (orange) or a good (green) alternative.
To make sure, our insights are not based on some unfaithful attributions, we also compare quantitative faithfulness for the and observe an average increase in the ABC metric by more than 10 % (see Sec. 5).
4.3 Why does the wind turbine produce less than expected?
Wind power is one of the pillars of decarbonizing energy systems around the world. Wind turbines are often placed in remote locations and need to be operated and monitored from a distance, using data from their Supervisory Control and Data Acquisition (SCADA) system. Effectively leveraging this data is an active area of research (Innes Murdo Black & Kolios, 2021), with the primary focus on detecting and diagnosing underperformance as the central challenge (Pandit et al., 2023). However, the detection of underperformance is always context-specific since the implicit question is: ‘underperformance relative to what?’ In the wind turbine case, it is the condition without the presence of a malfunction, given the context of prevailing ambient conditions.
We utilize data from a 2 MW wind turbine and a meteorological met-mast from an onshore wind farm on the Iberian peninsula  [^2] https://opendata.edp.com. SCADA data is available for two years and includes ambient conditions as well as technical turbine parameters as 10-minute averaged values (50,000 data points after pre-processing). We have trained a small fully-connected MLP to predict the turbine output from wind speed, air density, and turbulence intensity. The model achieves a competitive RMSE of less than 36 kW. Additionally, we have augmented the data with so-called yaw-misalignment losses. They occur when a turbine does not perfectly face the incoming wind direction, which reduces the effective area of the rotor. Detecting yaw-misalignment is part of ongoing research (Pandit et al., 2021; QU2, 2022) and attributing it by XAI methods has recently been proposed as an effective solution (Letzgus & Müller, 2024). For such an approach to work, we need our XAI methods to faithfully attribute the losses induced by yaw-misalignment to the respective feature (difference of nacelle and wind direction). In our setup, we can directly compare attributions with the respective ground-truth-losses.
Figure 6: Quantitative faithfulness when attributing yaw-misalignment losses to the respective feature with standard LRP (grey) and XpertAI-LRP (red) against the true losses (dashed line).
We trained five range experts for the different operational regions of the turbine. Figure 6 shows the comparison of attributing the yaw-misalignment induced losses to the respective yaw-feature with standard LRP and our XpertAI-LRP variant. We can observe that the naive LRP application attributes exhibit a systematic overestimation (larger negative values) of losses caused by incorporating phenomena from outside the respective operational regime. Our proposed novel attributions obtained from the range-experts, on the other hand, are on average much closer to the ground truth. For turbine operators, this directly translates to better operation and maintenance decisions and therefore highlights the benefit of using sub-manifold-specific explanations in industrial or engineering applications.
5 Quantitative Evaluation and Sanity Checks of XpertAI-faithfulness
After having shown some intriguing insights that were enabled by our XpertAI approach in the previous chapter, we now conduct a systematic evaluation of their faithfulness. We utilize the controlled environment of the previously introduced rMNIST dataset to investigate practically relevant questions of how to set up, train, and utilize our range experts. Furthermore, we add some sanity-check type evaluation to empirically validate Proposition 1 from Sec. 3 and get practically relevant insights for range-expert model selection. Details on the respective experiments can be found in Appendix C and some guidance on how to obtain faithful range experts in Appendix B.
5.1 Are XpertAI attributions faithful?
To answer this question properly, we first need to define how we measure faithfulness. In general, regression-specific evaluation of XAI attribution methods has not received much attention so far, (Hama et al., 2023) being a noteworthy exception. The authors propose a regression-specific flavor of the area under the curve in pixel-flipping (Bach et al., 2015; Samek et al., 2017), called the area between the curves (ABCs). The ABC is defined as the area between the model output when occluding a sample’s features in the order of attribution magnitudes and a straight line connecting $f(x)$ and $f(x^{\prime})$ (which corresponds to random sorting). Since sorting ascending and descending can result in asymmetrical curves, we sum over both areas (Blücher et al., 2024). We furthermore normalize by the distance between the sample and the baseline when averaging. Higher values of ABC are better.
Furthermore, the challenge of including context in attribution methods (Sec. 2.2) naturally extends to occlusion-based evaluation (ergo, what to occlude with?). Here again, the best way to ensure that we evaluate attributions within the relevant output range of function $f(x)$, where we account for context-specific (local) effects, is to occlude with a domain-specific counterfactual (Albini et al., 2022; Hama et al., 2023). Therefore, we sample conditional $x^{\prime}=D(x|y=\widetilde{y})$ from the available data set $D$ with which we then occlude and average the respective ABCs over multiple draws.
Table 1: Comparison of faithfulness for different attribution methods applied naively and within the XpertAI framework. Relative improvement of ABC over naive application.
dataset $LRP$ $IG$
rMNIST +42 % +7.2 %
WINE +23.6 % +10.6 %
FRIEDMAN +12.7 % +1.9 %
CALIFORNIA +1.8 % +9.7 %
DIABETES +2.6 % +4.4 %
Table 1 reports the ABC scores of our XpertAI approach relative to a naive application of LRP and IG on the previously introduced as well as several popular regression benchmarks (Pedregosa et al., 2011). For each of them, we trained three range experts and evaluated samples from the top range (see Appendix C). Overall, we see consistent improvements in ABC scores across all settings which means that our approach indeed can generate more faithful attributions with respect to a user-specific query. Note, that the advantage is significantly larger for LRP. There, the approach corresponds to a data-driven root-search strategy whereas naively, there is no such option. For IG we have already leveraged its inherent contextualization capability to some extent by utilizing the mean over all input samples as a starting point for the integration path. Our approach is still able to further refine the attributions towards a better contextualization.
Figure 7: Mean occlusion curves over all samples from the top bin of a six-expert-basis. When successively subtracting range-expert attributions from the original explanation and flipping pixels according to the remaining explanations, the flipping curves saturate in the proximity of the respective reference values.
5.2 (Diss-)aggregate XpertAI-attributions
From Proposition 1, we can in principle derive an alternative way to obtain disentangled and contextualized attributions with respect to $\widetilde{y}$. Instead of adding up the respective expert attributions, we subtract them from the original explanation in reverse order. Intuitively, only information relevant for higher-range bins should remain. We test this hypothesis empirically on the rMNIST dataset. We flip pixels to zero according to the order of the difference of attributions $\mathcal{E}(y,x)-\sum_{m}{\mathcal{E}(z_{m},x)}$. Intuitively, the more evidence associated with lower-range concepts we subtract, the more evidence for higher values should remain, and therefore the flipping curve should decrease more slowly. In the ideal case, the only information with positive attributions is the one relevant for values larger than $\widetilde{y}_{i}$ and the flipping curve should therefore remain around that value for as long as possible. When testing this empirically, we indeed see such a behavior (Figure 7). Note also, that the plateaus of the different range experts do not cluster around the digit-transitions (0.33 and 0.66). This means, that despite the strong global concept shifts present in the data, the range experts were able to capture more subtle, local effects that guide $f(x)$ in the context of the respective reference values.
Table 2: Results for pixel flipping experiments for regression MNIST. Results within ranges: sample-flipping baseline pairs are within one expert range. ABC values are normalized by flipping distance, high values are better.
##\## $LRP$ $LRP$ $IG$ $IG$
$experts$ $Naive$ $XpertAI$ $Naive$ $XpertAI$
3 0.40 0.56$\pm$ 0.02 0.93 1.00 $\pm$ 0.05
5 0.47 0.70$\pm$ 0.01 1.16 1.20 $\pm$ 0.01
6 0.46 0.75$\pm$ 0.01 1.22 1.23$\pm$ 0.01
9 0.49 0.78$\pm$ 0.01 1.32 1.36$\pm$ 0.01
5.3 How many range experts?
One practically very relevant question is, how many range experts to train, which includes the choice of their respective ranges ($\tau$). Conceptually, the method should work best if every distinct sub-region of the output is covered by at least one range expert. In practice, these can either be domain-informed and therefore known apriori, or inferred by analyzing activation patterns (from an activation vs. $f(x)$ scatter plot, for example). In the context of our rMNIST case, selecting one range expert for each digit range, therefore three range experts in total appears to be the most intuitive choice. In practice, however, we might not know where exactly these boundaries lay. Therefore, we compare settings with three, five, six, and nine equally spread range experts.
In Table 2 we see that our approach improved the ABC score across all settings. Also, we can see that LRP benefits in particular from adding extra range experts while IG results are more consistent across the number of experts. Note, that this also holds if the expert ranges are not aligned with known sub-concept ranges (as is the case for 5 equally distributed experts). In practice this means that the limit for the number of experts depends on the specific problem, computational considerations as well as the resolution of the available data.
6 Discussion and Conclusions
In this paper, we have proposed the XpertAI framework to achieve contextualized and disentangled attributions when explaining regression models. Inspired by the MoE approach, the framework divides the data into sub-manifolds, each of which corresponding to a certain predicted output range. Such a division is achieved by building a collection of range experts, which we equip with explainability. It enables for the first time a disentanglement along the output of the prediction strategy and the resolution of specific user-defined queries.
Empirically, we find that our XpertAI framework is able to distill locally relevant explanations from highly aggregated global standard attributions, as demonstrated by a number of quantitative experiments based on occlusion tests. Explanations associated with each expert range can be precomputed, in a way that exact user queries can be answered very quickly as a linear combination of the precomputed explanations.
Our approach can be interpreted within the framework of virtual layers, which has been instrumental in achieving various forms of explanation disentanglement. Furthermore, our approach provides an alternative to the more common approach of extracting reference points or counterfactuals and bypasses some of the challenges associated with reference points, such as their multiplicity and the need to search for them. Also, our approach differs from self-interpretable generalized additive models, by remaining applicable to a broad range of ML models, including deep neural networks.
While we have demonstrated that our method can work alongside various explanation techniques, in particular, gradient-based techniques such as Integrated Gradients, or propagation-based techniques such as LRP, it is necessary for the latter to structurally disentangle the range experts. While we have proposed for this later step a surrogate modeling approach, these surrogates need to be carefully trained and regularized to maintain the original model’s prediction output as well as its prediction strategy. Hybrid approaches, with the top layers handled by perturbation-based techniques and the lower layers with propagation, may eliminate the need for structural disentanglement while at the same time retaining high accuracy and computational efficiency.
Overall, our work has highlighted the need to precisely formulate “what to explain” (the explanandum) and proposed a practical and flexible solution in the context of regression. The MoE idea our method relies on, however, is more general, and our framework could be extended in the future to other decomposition of the predicted output, e.g. for structured output tasks such as time series prediction.
Impact statement: This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
Appendix A Proof of Propositions 1 and 2
Proposition 1 stating the conservation property of the proposed query explanation can be demonstrated through the chain of equations:
$\textstyle\sum_{i}\mathcal{E}(q,x)_{i}$ $=\textstyle\sum_{i}\sum_{m}w_{m}\mathcal{E}(z_{m},x)_{i}$ (10)
$=\textstyle\sum_{m}w_{m}\sum_{i}\mathcal{E}(z_{m},x)_{i}$ (11)
$=\textstyle\sum_{m}w_{m}z_{m}$ (12)
$=q$ (13)
where in (10), we have injected the expression of the explanation in (9). From (10) to (11) we have permuted the sums. From (11) to (12), we have used the conservation property of the explanation of $z_{m}$. From (12) to (13) we have identified the weighted sum as being the query. Likewise, for Proposition 2, if some feature $i$ satisfies $\forall_{m}:\mathcal{E}(z_{m},x)_{i}=0$, then
$\mathcal{E}(q,x)_{i}$ $=\textstyle\sum_{m}w_{m}\mathcal{E}(z_{m},x)_{i}$ (14)
$=\textstyle\sum_{m}w_{m}\cdot 0$ (15)
$=\textstyle 0$ (16)
Appendix B How to train and select good range experts?
In practice, we need to train and select appropriate range experts for the XpertAI approach to enhance contextualization as reported above. This process may vary based on the respective XAI attribution method being employed. For occlusion- and gradient-integration-based methods, which do not require additional structural disentanglement (see Section 3.3), a simple shift-and-clip strategy is sufficient. For propagation-based methods, however, we need to learn the surrogate $a\mapsto(z_{m})_{m}$, as described earlier. Here, we want to highlight the need for appropriate regularization to avoid overfitting, which in the case of range experts would result in unfaithful model attributions. Consequently, we want to choose the least complex range expert, that can sufficiently learn the respective mapping.
In case the latent representation $a$ is already adequately disentangled, it is sufficient to fit a linear range expert (without bias term). We have observed this to work well for some of our low-dimensional benchmark datasets. Otherwise, we need to gradually increase range-expert complexity (adding neurons and reducing L2-regularization) until the mapping is learned sufficiently. Moreover, we have observed that instead of additional layers, (copying and) fine-tuning the top layer(s) on the range-expert targets $z_{m}$ with small learning rates is a good strategy since it ensures the solution lays in relative proximity to the original model. If a new layer is added, initializing the weights with a projection to the latent principal components conditioned on the respective output range ($PCA(X|z_{m})$) was found to speed up training and ensure good results. Furthermore, the Shapley-style data augmentation (cf. Section 3) is another crucial ingredient to prevent our experts from adhering to spurious correlations (that our original models did not use). This can be conveniently implemented with the help of a dropout layer on the surrogate input $\boldsymbol{a}$. Lastly, we can enforce the saturation of range experts outside their area of expertise by adding an explicit combination of ReLU functions that clip $s_{m}$ to the desired range. These measures together ensure faithful and computationally efficient range experts.
Appendix C Details on Evaluation (Sec. 4 and 5)
Figure 8: Details on datasets, models and range experts for the respective benchmark datasets
Figure 8: Details on datasets, models and range experts for the respective benchmark datasets
Figure 9: Mean attributions over different cohorts of samples (columns) and basis functions (rows). Equivalent plot to Fig. 4 but for six range expert basis functions.
Figure 10: Overview model performance for the regression benchmarks presented in Sec. 5.Left: original label distribution colour-coded by range expert. Centre: original label vs. original model output. Right: original model output vs. output of the respective range experts. Note, that the results shown are from range experts retrained to obtain a more disentangled representation in connection with LRP.
Figure 11: Overview of model performance on the rMNIST problem for 3, 6 and 9 range experts (top to bottom). Left: original label distribution colour-coded by range expert. Centre: original label vs. original model output. Right: original model output vs. output of the respective range experts. Note, that the results shown are from range experts retrained to obtain a more disentangled representation in connection with LRP.
Appendix D Augmenting wind turbine SCADA data with yaw-misalignment
We randomly add yaw misalignment of up to $15^{\circ}$ to our data SCADA set, and adjust the respective targets (turbine output) with a yaw misalignment factor $c_{ymis,i}=cos^{3}(\Delta_{yaw})$, if $v_{w,i}<v_{w,rated}$. This approximation can be easily derived from static flow equations and geometric considerations, for more details on how yaw misalignment affects turbine output see (Howland et al., 2020). After training and evaluation of the model on the augmented data, we can compare the magnitude of attributions to the ground truth.